# Feature Hedging: Correlated Features Break Narrow Sparse Autoencoders

### v1

@@ 3.4 Anti-correlated features
Next, we reverse the conditional probabilities of $f_{2}$ so that $P(f_{2}|f_{1})=0.1$ and $P(f_{2}|\neg f_{1})=0.2$. Now $f_{2}$ is more likely to fire on its own than it is to fire along with $f_{1}$. A plot of the cosine similarity of the SAE with the true features is shown in Figure 3(b).
@@
Now the SAE latent has actually merged a negative component of $f_{2}$ into its single latent instead of a positive component. Furthermore, increasing L1 penalty does nothing to solve this, as the negative component of hedging in the encoder does not increase L0 of the SAE. If we use a full-width SAE, we again see the SAE learns the true features despite the correlation (see Appendix A.1).
3.5 Hedging is caused by reconstruction loss: curves for single-latent SAEs
@@ 4 Quantifying hedging in LLM SAEs
While we have demonstrated hedging in a synthetic setting, it remains a question how much hedging occurs in LLM SAEs. We next study the effect of adding new latents to an existing SAE. Based on our understanding of hedging in toy models, we expect that when a new latent is added to an SAE, this should pull the component of the new feature out of existing SAE latents. Thus if hedging occurs, the change in existing latents after a new latent is added should project onto that new latent. If hedging did not exist, then adding a new latent should not have any effect on existing latents.
Hedging affects the encoder and decoder of the SAE symmetrically, so we should be able to detect hedging in either the encoder or decoder. We look at the decoder to distinguish hedging from absorption, as absorption affects the encoder. Under feature absorption, if a newly added latent is a child feature of an existing latent, then the encoder for the parent latent adds a negative component of the new child latent to avoid firing when the child is active, but the parent decoder remains unchanged. This corresponds to adding a new latent to Figure 1(a) and arriving at Figure 1(b). Thus, any change to existing decoder latents cannot be attributed to absorption and must be due to hedging.
We expect that even if there were no hedging at all, simply due to noise, existing SAE decoder latents may undergo a change that has some small projection onto new added latents. We want to make sure that anything we quantify as hedging must be larger than what we would expect from random noise.
@@
Taking this into account, we define a metric called hedging degree, $h$. We take an existing SAE $s_{0}$ with $L$ latents and add $N$ new latents to the SAE. After adding these latents, we continue training the SAE and arrive at a new SAE, $s_{1}$, with $L+N$ latents. We also continue training $s_{0}$ on the same tokens that we train $s_{1}$ on to ensure that any difference between $s_{0}$ and $s_{1}$ is due only to the newly added latents. $W^{0}_{\text{dec}}$ refers to the new decoder of $s_{0}$, and $W^{1}_{\text{dec}}$ refers to the decoder of $s_{1}$. We define the difference in the original $L$ latents between $s_{0}$ and $s_{1}$ as:
@@
where $W^{1}_{\text{dec}}[L:L+N]$ refers to the newly added decoder latents. $W_{\text{rand}}[0:N]$ refers to a decoder consisting of $N$ randomly initialized latents. All decoders are normalized to have latents of unit norm. We define the projection of a vector $v$ onto a subspace spanned by $W$ as:
@@
The hedging degree $h$ is then defined as:
$h=\frac{1}{L}\sum^{L}_{i}\underbrace{\|\text{Proj}(\delta_{L}[i],W^{1}_{\text{dec}}[L:L+N])\|}_{\text{Projection of $\delta_{L}$ onto N new latents}}-\underbrace{\|\text{Proj}(\delta_{L}[i],W_{\text{rand}}[0:N])\|}_{\text{Projection of $\delta_{L}$ onto N random latents}}$ (7)
Any value of $h>0$ corresponds to hedging above what we would expect from random noise, as $h$ subtracts the projection along $N$ randomly initialized latents as part of the computation.
@@ 4.1 Results
We experiment with SAEs trained on Gemma-2-2b [21], as this model is commonly used for SAE research due to the thoroughness of the Gemma Scope suite of SAEs [15], as well as Llama-3.2-1b [7] to validate results on another LLM. All SAEs are trained first on 250M tokens of the Pile uncopyrighted [10]. After adding $N=64$ latents, we continue training for another 250M tokens. The version of the SAE without latents added is also trained for another 250M tokens, so each SAE is trained for 500M tokens total. The pair of extended and non-extended SAEs is used to calculate hedging degree. SAE training details are in Appendix A.2.
@@
It also appears that BatchTopK SAEs have more hedging than L1 SAEs. This may be due to L1 loss reducing hedging from positively correlated features, as we saw in Section 3.3.
5 Case study: adding a new latent to an existing SAE
We next explore how hedging affects a real SAE. We trained a L1 SAE on Gemma-2-2b layer 12 with width 8192 for 250M tokens on the Pile [10], then add a new latent to the SAE, and continue training both the original SAE and the extended SAE for another 250M tokens.
0.3/css/bootstrap.min.
/bootstrap.min.css_integrity="sha3"
/bootstrap.min.css">____link
>____<link_rel="stylesheet"_type="
8"/>< link_rel=stylesheet_href="..//doc
png"><link_rel="manifest"_href="
(b) Latent 3094, which had the largest negative $\delta$-projection after adding latent 8192. This latent tracks ‘‘rel’’ in HTML, used for CSS scripts in HTML.
Figure 7: Sample top activating examples for case study latents.
@@ 6 Balancing hedging and absorption in matryoshka SAEs
Matryoshka SAEs [4] combat absorption with nested SAE loss prefixes. Each level acts like a small SAE, and is forced to reconstruct the input on its own. This forces the SAE to learn more general concepts in earlier levels, and makes it difficult for the SAE to make holes in the recall of parent latents for absorption, as this would hurt the reconstruction of earlier matryoshka levels.
However, since early matryoshka levels are effectively narrow SAEs, they suffer from feature hedging. As we saw in Section 4.1, the more narrow an SAE is, the more the severe the feature hedging. Matryoshka SAEs thus solve feature absorption at the expense of exacerbating feature hedging.
@@
We demonstrate this balancing in a toy model of hierarchical features. The toy model has 4 features, with feature 1 being the parent feature and features 2-4 being children (features 2-4 can only fire if feature 1 is also firing). Feature 1 fires with probability 0.25, and each child feature fires with probability 0.15 if feature 1 is firing. We train a matryoshka SAE with a single inner level consisting of only latent 1 with balance coefficient $\beta$. For more details on this toy setup, see Appendix A.5.
(a) Matryoshka SAE with detached loss (equivalent a matryoshka SAE with $\beta=\infty$). Hedging adds positive components of the child features 2-4 to the encoder of latent 1.
@@
Figure 8: Balancing hedging and absorption in a toy model of hierarchical features. Child features 2-4 only fire if parent feature 1 fires. The matryoshka SAE has a single inner level with 1 latent.
@@
We train 10 seeds for each multiplier and show results in Figure 9 for absorption rate, targeted probe pertubation (TPP), Spurious Concept Removal (SCR), K-sparse probing, and feature-splitting metrics from SAEBench [14], and k=1 sparse probing results [12] for a Parts of Speech (POS) dataset we created using Treebank POS tagged sentences [16]. We add a POS dataset for probing since POS are very general concepts, and should be learned in the earliest levels of a matryoshka SAE.
For TPP, feature splitting, and sparse probing, using a compound multiplier of around 0.75 achieves better results than either a standard matryoshka SAE or a standard (non-matryoshka) SAE, providing evidence that balancing matryoshka losses can improve the performance. Using a multiplier of 0.75 still scores well on the absorption metric as well. Strangely, the SCR metric appears to perform better at higher multipliers. SCR and TPP should measure the same thing, so we do not fully understand the discrepancy between these metrics. We provide further results and more details in Appendix A.7.
While balancing each $\beta_{m}$ can improve performance on most metrics, we do not expect this to perfectly solve absorption and hedging. We show in Appendix A.6 that balancing all hedging and absorption with a single $\beta_{m}$ is not always possible. We thus expect that it may be possible to further improve SAE performance by learning different balancing coefficients per latent, but this is left to future work.
@@ 7 Related work
Other work has highlighted theoretical problems with SAEs. Till [23] investigated a problem where SAEs may increase sparsity by inventing features. For instance, an SAE may fabricate a “red triangle” feature in addition to “red” and “triangle” features. Templeton et al. [22] dicuss the problem of feature splitting, where an SAE may not learn features at a desired level of specificity. Engels et al. [8] investigates SAE errors and finds that SAE error may be pathological and non-linear. Engels et al. [9] further shows that there are features that cannot be expressed as a simple linear direction, and thus SAEs may struggle to represent these features. Wu et al. [24] and Kantamneni et al. [13] both investigate the empirical performance of SAEs and find that SAEs underperform baselines.
@@ 8 Discussion
SAEs remain a promising technique for decomposing the residual stream of LLMs in an unsupervised manner. However, given recent work showing that SAEs underperform relative to baselines [24, 13], it is imperative that we understand the reasons for this underperformance so they can be addressed.
@@ 9 Limitations
Appendix A Technical Appendices and Supplementary Material
A.1 Full-width SAE toy model results
@@
We extend the discussion of single-latent SAEs to explore what happens if the SAE has two latents, the same number of latents as the number of true features. We use the same toy model as in Section 3.3 for the positive correlation case, and the same toy model as in Section 3.4 for the anti-correlated case. We use L1 penalty of 1e-3 for the positive correlation case, the same as the L1 penalty that caused hedging in single-latent SAEs.
@@
A.2 Training details for LLM SAEs
All SAEs are trained on the Pile uncopyrighted [10], using a batch size of 4096 activations and context length of 1024 tokens. For L1 SAEs, we use a linear L1 warm-up of 10k steps. SAEs are trained on a single 80gb Nvidia H100 GPU. Model weights are loaded in fp32 precisions, but autocast to bfloat16 during training. We initialize the SAE so that the encoder and decoder are identical, where each latent has norm 0.1, following the procedure described in [18]. All L1 SAEs are trained with learning rate 7e-5, and BatchTopK SAEs are trained with learning rate 3e-4. SAEs are trained using SAELens [1].

### v2

@@ 3.4 Anti-correlated features
So far we have only seen the effect of positive correlation between features. We next change our toy model so $f_{4}$ is more likely to fire if $f_{3}$ does not fire. We set $p_{4}=0.1$ if $f_{3}$ fires, but $p_{4}=0.2$ if $f_{3}$ does not fire. Results are shown in Figure 5.
Figure 5: SAEs trained on a toy model with negative (anti) correlation between features $f_{3}$ and $f_{4}$. When the SAE is too narrow to represent $f_{4}$ (left), we latent 3 mixes a negative component of $f_{4}$. When the SAE is wide enough to contain both $f_{3}$ and $f_{4}$ (right), the SAE learns correct features.
We now see that $l_{4}$ is mixing in a negative component of $f_{3}$. This demonstrates that the correlation is the cause of the hedging: flipping the sign of the correlation flips the sign of the hedging.
The implications of this for SAE performance are quite dire. While it is already bad for positively correlated features to become hedged (e.g. “sunshine” and “summertime”), at least the mixed features have some relation to each other. For anti-correlated features, this could look like a latent for “chemical molecules” having a negative component of the “Darth Vader” feature mixed in, since chemical modules and Darth Vader are highly anti-correlated. Worse, it is not even clear if the inverse of the “Darth Vader” direction is meaningful in the model at all, or is just noise. We further expect that there are many more negative correlations than positive correlations in language, e.g. a negative component every word in every non-English language may be mixed into every latent tracking an English word. These negative correlations likely introduce what looks like a lot of random noise into SAE latents, and this can only harm performance and interpretability.
3.5 Studying hedging in single-latent SAEs
We next investigate hedging in the simplest possible toy SAE setting: an SAE with a single latent. We use a model with two true features $f_{1}$ and $f_{2}$ ($N=2,D=50$). Each feature fires with magnitude 1.0. Unless otherwise specified, $f_{1}$ fires with probability 0.25, and $f_{2}$ fires with probability 0.2. We use SAELens (Bloom et al., 2024) to train a single-latent SAE on these activations.
3.5.1 Fully independent features
Figure 6: True features, and SAE decoder latent and $b_{\text{dec}}$ for single-latent SAE and a toy model with two true features. When the features fire independently, there is no hedging seen in the SAE latent. When any correlation is present, the SAE latent shows clear hedging.
We first study the case when $f_{1}$ and $f_{2}$ fire independently. We find that the SAE correctly represents $f_{1}$ without any interference from $f_{2}$. However, the decoder bias has incorrectly learned to represent the direction of $f_{2}$, but with magnitude 0.2, equal to the probability of $f_{2}$ firing. The single SAE latent, SAE bias term, and true features are shown in Figure 6a.
We consistently find this pattern of the decoder bias merging in positive components of features not tracked by their own latent. In this sense, the decoder bias can be thought of as an always-on latent, and thus is thus also susceptible to hedging.
3.5.2 Hierarchical features
Next, we investigate what happens if $f_{1}$ and $f_{2}$ are in a hierarchy, so $f_{2}$ can only fire if $f_{1}$ fires, but $f_{1}$ can still fire on its own ($f_{2}\implies f_{1}$). We adjust the firing probability of $f_{2}$ so that $P(f_{2}|f_{1})=0.2$, and $P(f_{2}|\neg f_{1})=0$ (thus, $P(f_{2})=0.05$). In a two-latent SAE this setup would cause feature absorption. We plot the SAE latent, decoder bias, and true features in Figure 6b.
Here we clearly see feature hedging. The single SAE latent has now merged in a component of $f_{2}$ into its single latent, so it is now a mixture of $f_{1}$ and $f_{2}$. This merging of features reduces the MSE loss of the SAE despite being a degenerate solution.
3.5.3 Positively correlated features
Next, we change our setup so that $P(f_{2}|\neg f_{1})=0.1$ instead of 0. We still keep $P(f_{2}|f_{1})=0.2$, so that $f_{2}$ is more likely to fire if $f_{1}$ fires, but it can still fire on its own as well. The features are now merely correlated rather than following a strict hierarchy. Results are shown in Figure 6c.
We still see hedging in the SAE latent, but less than with full hierarchical features. However, if the L1 penalty is high enough and the level of correlation is low enough, then the SAE can still learn the correct features, as positive hedging increases the L0 of the SAE slightly relative to learning just $f_{1}$. We show the resulting SAE latent and features with high L1 penalty in Figure 7(b). Interestingly, we now see that the hedging has moved more apparently into the decoder bias instead. If we use a full-width SAE, the SAE learns the true features despite the correlation (Appendix A.2).
3.5.4 Anti-correlated features
@@
Next, we reverse the conditional probabilities of $f_{2}$ so that $P(f_{2}|f_{1})=0.1$ and $P(f_{2}|\neg f_{1})=0.2$. Now $f_{2}$ is more likely to fire on its own than it is to fire along with $f_{1}$, making these feaures anti-correlated. Results are shown in Figure 6d.
Now the SAE latent has actually merged a negative component of $f_{2}$ into its single latent instead of a positive component. How does this work? We see that the decoder bias, $b_{\text{dec}}$, has a larger component of $f_{2}$ than in the positive correlation case. The SAE is using the decoder bias to include a “default” value for $f_{2}$, and then when $f_{1}$ fires, the SAE latent’s negative component of $f_{2}$ acts to reduces the amount of $f_{2}$ present in the reconstruction. The SAE is abusing the correlation to adjust its guess of the amount of $f_{2}$ that should be output despite not having a dedicted latent for $f_{2}$: if $f_{1}$ is active, then the likelihood that $f_{2}$ is active decreases, and the SAE likewise reduces the amount of $f_{2}$ that is output.
@@
Increasing L1 penalty cannot solve this, as the negative component of hedging in the encoder does not increase L0 of the SAE. If we use a full-width SAE, we again see the SAE learns the true features despite the correlation (see Appendix A.2).
3.5.5 Hedging is a function of feature correlation
Next, we explore the effect of feature correlation on the amount of hedging in our single-latent, two feature setting. We set $P(f_{1})=0.45$ and $P(f_{2})=0.25$, but change the correlation between these features, $\rho$, to range from 0.5-0.5 to $0.5$. We then calculate the cosine similarity of the SAE decoder latent, $l$, with $f_{2}$. We furthermore initialize the single SAE latent to match $f_{1}$, so that any deviation from this must be caused by gradient pressure rather than simply being an unfortunate local minimum. If there is no hedging occurring, then $\cos(l,f_{2})=0$, as we saw in Figure 6a. Results are shown in Figure 7(a).
(a) Hedging amount ($\cos(l,f_{2})$) vs correlation between $f_{1}$ and $f_{2}$. The degree to which $l$ mixes in $f_{2}$ is a clear function of the amount of correlation between features.
(b) High L1 penalty can reduce hedging caused by positive correlations.
Figure 7: Hedging as a function of feature correlation, and effect of L1 penalty on positive hedging.
As expected, the amount of hedging directly tracks the amount of correlation. The hedging also matches the sign of the correlation as well, with negative correlation resulting in a negative component of $f_{2}$ being mixed into $l$, and positive correlation resulting in a positive component of $f_{2}$ being mixed into $l$.
@@ 4 Quantifying hedging in LLM SAEs
While we have demonstrated hedging in a synthetic setting, it remains a question how much hedging occurs in LLM SAEs. Based on our understanding of hedging in toy models, we expect that when a new latent is added to an SAE, this should “pull out” the component of the new feature from existing SAE latents, where it was previously hedged. Thus if hedging occurs, the change in existing latents after a new latent is added should project onto that new latent. If hedging did not exist, then adding a new latent should not have any effect on existing latents.
Parent latents are learned before child latents
A key assumption in Matryoshka SAEs is that if latents exist in a hierarchy, and the SAE is too narrow to represent both the parent and child, the SAE will learn the parent first. We feel this assumption is reasonable since parent latents, by definition, fire more frequently than child latents, so the SAE is incentivized to learn them first.
This insight allows us to differentiate hedging from absorption. Under absorption, if a newly added latent is a child feature of an existing latent, then the encoder for the parent latent adds a negative component of the child latent to avoid firing when the child is active, but the parent decoder latent remains unchanged. This corresponds to adding $l_{2}$ to Figure 1(a) and arriving at Figure 1(b). The decoder of $l_{1}$ (the parent) remains identical to before $l_{2}$ is added, except the hedging from $f_{2}$ is removed. Thus, changes to existing decoder latents cannot be absorption and must be due to hedging.
@@
Taking this into account, we define a metric called hedging degree, $h$. We take an existing SAE $s_{0}$ with $L$ latents and add $N$ new latents to the SAE. After adding these latents, we continue training the SAE and arrive at a new SAE, $s_{1}$, with $L+N$ latents. We also continue training $s_{0}$ on the same tokens that we train $s_{1}$ on to ensure that any difference between $s_{0}$ and $s_{1}$ is due only to the newly added latents. $W^{0}_{\text{dec}}$ refers to the new decoder of $s_{0}$, and $W^{1}_{\text{dec}}$ refers to the decoder of $s_{1}$. $W_{dec}$ is normalized so each latent has unit norm. We define the difference in the original $L$ latents between $s_{0}$ and $s_{1}$ as:
@@
where $W^{1}_{\text{dec}}[L:L+N]$ refers to the newly added decoder latents. $W_{\text{rand}}[0:N]$ refers to a decoder consisting of $N$ randomly initialized unit-norm latents. All decoders are normalized to have latents of unit norm. We define the projection of a vector $v$ onto a subspace spanned by $W$ as:
@@
We expect that even if there were no hedging at all, simply due to noise, existing SAE decoder latents may undergo a change that has some small projection onto new added latents. We want to make sure that anything we quantify as hedging must be larger than what we would expect from random noise. Taking this into account, the hedging degree $h$ is then defined as:
[…]
Any value of $h>0$ corresponds to hedging above what we would expect from random noise, as $h$ subtracts the projection along $N$ randomly initialized unit-norm latents as part of the computation.
@@ 4.1 Results
We experiment with SAEs trained on Gemma-2-2b (Team et al., 2024), as this model is commonly used for SAE research due to the thoroughness of the Gemma Scope suite of SAEs (Lieberum et al., 2024), as well as Llama-3.2-1b (Dubey et al., 2024) to validate results on another LLM. All SAEs are trained first on 250M tokens of the Pile uncopyrighted (Gao et al., 2020). After adding $N=64$ latents, we continue training for another 250M tokens. The version of the SAE without latents added is also trained for another 250M tokens, so each SAE is trained for 500M tokens total. The pair of extended and non-extended SAEs is used to calculate hedging degree. SAE training details are in Appendix A.3.
@@
It also appears that BatchTopK SAEs have more hedging than L1 SAEs. We suspect that L1 loss can reduce hedging from positively correlated features, as we saw in Section 3.5.3.
We further validate hedging in LLM SAEs via a case-study of adding a new latent to an SAE trained on Gemma-2-2b in Appendix A.5.
@@ 5 Balancing hedging and absorption in matryoshka SAEs
Matryoshka SAEs (Bussmann et al., 2025) combat absorption with nested SAE loss prefixes. Each level acts like a small SAE, and is forced to reconstruct the input on its own. This forces the SAE to learn more general concepts in earlier levels, and makes it difficult for the SAE to make holes in the recall of parent latents for absorption, as this would hurt the reconstruction of earlier levels.
However, since early matryoshka levels are effectively narrow SAEs, they suffer from feature hedging. As we saw in Section 4.1, the more narrow an SAE is, the worse the hedging. Matryoshka SAEs thus solve feature absorption at the expense of exacerbating feature hedging.
@@
We demonstrate this balancing in a toy model of hierarchical features. The toy model has 4 features, with feature 1 being the parent feature and features 2-4 being children (features 2-4 can only fire if feature 1 is also firing). Feature 1 fires with probability 0.25, and each child feature fires with probability 0.15 if feature 1 is firing. We train a matryoshka SAE with a single inner level consisting of only latent 1 with balance coefficient $\beta$ (Since there is only one inner level, we always set the outer level coefficient to 1). For more details on this toy setup, see Appendix A.7.
(a) Matryoshka SAE with detached loss (equivalent to a matryoshka SAE with $\beta=\infty$). Hedging adds positive components of the child features 2-4 to the encoder of latent 1.
@@
Figure 9: Balancing hedging and absorption in a toy model of hierarchical features. Child features 2-4 only fire if parent feature 1 fires. The matryoshka SAE has a single inner level with 1 latent, represented by a black box around latent 1.
@@
We train 10 seeds for each multiplier and show results in Figure 10 for absorption rate, targeted probe pertubation (TPP), Spurious Concept Removal (SCR), K-sparse probing, and feature-splitting metrics from SAEBench (Karvonen et al., 2025), and k=1 sparse probing results (Gurnee et al., 2023) for a Parts of Speech (POS) dataset we created using Treebank POS tagged sentences (Marcus et al., 1993). We add a POS dataset for probing since POS are very general concepts, and should be learned in the earliest levels of a matryoshka SAE.
For TPP, feature splitting, and sparse probing, using a compound multiplier of around 0.75 achieves better results than either a standard matryoshka SAE or a standard (non-matryoshka) SAE, providing evidence that balancing matryoshka losses can improve the performance. Using a multiplier of 0.75 still scores well on the absorption metric as well. Strangely, SCR appears to perform better at higher multipliers. However, SCR is also the noisiest metric, and the noise is higher at high multipliers, so it could be that hedging increases the noise of the SCR metric but does not fully break it. We provide further results and more details in Appendix A.9.
While balancing each $\beta_{m}$ can improve performance on most metrics, we do not expect this to perfectly solve absorption and hedging. We show in Appendix A.8 that balancing all hedging and absorption with a single $\beta_{m}$ is not always possible. We expect it may be possible to further improve performance by learning different balancing coefficients per latent, but this is left to future work.
@@ 6 Related work
Other work has highlighted theoretical problems with SAEs. Till (2024) investigated a problem where SAEs may increase sparsity by inventing features. For instance, an SAE may fabricate a “red triangle” feature in addition to “red” and “triangle” features. Templeton et al. (2024) dicuss the problem of feature splitting, where an SAE may not learn features at a desired level of specificity. Engels et al. (2024) investigates SAE errors and finds that SAE error may be pathological and non-linear. Engels et al. (2025) further shows that there are features that cannot be expressed as a simple linear direction, and thus SAEs may struggle to represent these features. Wu et al. (2025) and Kantamneni et al. (2025) both investigate the empirical performance of SAEs and find that SAEs underperform baselines.
@@ 7 Discussion
SAEs remain a promising technique for decomposing the residual stream of LLMs in an unsupervised manner. However, given recent work showing that SAEs underperform relative to baselines (Wu et al., 2025; Kantamneni et al., 2025), it is imperative that we understand the reasons for this underperformance so they can be addressed.
@@ Acknowledgments
Appendix A Appendix
A.1 Hedging is caused by reconstruction loss: curves for single-latent SAEs
A.2 Full-width SAE toy model results
@@
We extend the discussion of single-latent SAEs to explore what happens if the SAE has two latents, the same number of latents as the number of true features. We use the same toy model as in Section 3.5.3 for the positive correlation case, and the same toy model as in Section 3.5.4 for the anti-correlated case. We use L1 penalty of 1e-3 for the positive correlation case, the same as the L1 penalty that caused hedging in single-latent SAEs.
@@
A.3 Training details for LLM SAEs
All SAEs are trained on the Pile uncopyrighted (Gao et al., 2020), using a batch size of 4096 activations and context length of 1024 tokens. For L1 SAEs, we use a linear L1 warm-up of 10k steps. SAEs are trained on a single 80gb Nvidia H100 GPU. Model weights are loaded in fp32 precisions, but autocast to bfloat16 during training. We initialize the SAE so that the encoder and decoder are identical, where each latent has norm 0.1, following the procedure described in (Olah et al., 2024). All L1 SAEs are trained with learning rate 7e-5, and BatchTopK SAEs are trained with learning rate 3e-4. SAEs are trained using SAELens (Bloom et al., 2024).